# Aerodynamics and Aeroacoustics of da Vinci’s Aerial Screw

### v1

@@ Appendix A Grid Convergence
We use the case with the highest Reynolds number of 8000 to verify grid convergence where a medium mesh contains 13.8 million grid points and a fine mesh with 32 million grid points is used. The region around the rotor is placed in a uniform fine mesh with a gradually expanding mesh used away from the rotor body. In the plane of the rotation, a $2.4R_{b}\times 2.4R_{b}$ region centered around the rotor is finely resolved while in the vertical axis, a length of $2R_{b}$ surrounding the screw is finely resolved (figure 2).
The lift coefficient is shown for the case with Reynolds number of 8000 in figure 9a and the error between the RMS value of the lift coefficient between these two mesh is 1.04%, showing that the results presented here are grid independent. The vortex field is also shown for the medium mesh (figure 9b) and the fine mesh (figure 9c) case with the vortices shown using the iso-surface of $Q$[28] defined as $\frac{1}{2}(||\bm{\Omega}||^{2}-||{\bf S}||^{2})$, where ${\bf S}$ and $\bm{\Omega}$ are the symmetric and anti-symmetric component of the velocity gradient tensor. We observe that the flow field looks visually similar on both the meshes indicating that not just the force but also the vortex structure are well converged. All the results for Reynolds 8000 case employ fine mesh while those with Reynolds number of 4000 and 2000 employ medium size mesh.

### v2

@@ V Acknowledgments
VI Author Declarations
The authors have no conflicts to disclose
@@ Appendix A Grid Convergence
We use the case with the Reynolds number of 8000 to verify grid convergence where a medium mesh contains 13.8 million grid points and a fine mesh with 32 million grid points is used. The region around the rotor is placed in a uniform fine mesh with a gradually expanding mesh used away from the rotor body. In the plane of the rotation, a $2.4R_{b}\times 2.4R_{b}$ region centered around the rotor is finely resolved while in the vertical axis, a length of $2R_{b}$ surrounding the screw is finely resolved (figure 2).
@@
The lift coefficient is shown for the case with Reynolds number of 8000 in figure 10a and the error between the RMS value of the lift coefficient between these two mesh is 1.04%, showing that the results presented here are grid independent. The vortex field is also shown for the medium mesh (figure 10b) and the fine mesh (figure 10c) case with the vortices shown using the iso-surface of $Q$. We observe that the flow field looks visually similar on both the meshes indicating that not just the force but also the vortex structure are well converged. All the results for Reynolds 8000 and 16000 case employ fine mesh while those with Reynolds number of 4000 and 2000 employ medium size mesh.
The rotating reference frame implementation (equation 2) is validated by simulating the aerial screw at Reynolds number of 2000 case till $t/T$ = 1.25 in both absolute and non-inertial rotating reference frame and the lift coefficient is shown for both the case in figure 11.
Figure 11: The lift coefficient is shown for the aerial screw at Reynolds number of 2000 and is simulated both in absolute frame and non-inertial rotating reference frame.